OpenStudy (anonymous):

Find the value of the expression of "4-2x" when x= (-2)

OpenStudy (anonymous):

4-2(-2)=?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

nope: -2(-2) =4

OpenStudy (anonymous):

so, 4-2(-2) = 4+4=

OpenStudy (anonymous):

You can do it rannsan!!!

OpenStudy (anonymous):

I got 8

OpenStudy (anonymous):

Good job!